OpenStudy (anonymous):

In October, a hardware store purchased snow shovels for $8 each. The original markup was 50% based on the selling price. In December, the store had a sale and marked down the shovels by 20%. By January 1, the sale was over and the shovels were marked up 15%. In March, the store held a clearance sale of all winter tools by having a final markdown of 60%. What was the final selling price of the shovels? (Round to the nearest cent.)

OpenStudy (anonymous):

I think the big issue with the second problem is the fact that the first markup is based on selling price. So, to calculate the selling price, you need to solve the equation below for y, where y represents the selling price. .5 = (y-8)/y You do this by multiplying both sides by y, to get .5y = y-8. Then subtracting y from both sides to get -.5y = -8. Then dividing both sides by -.5. This gives you the initial selling price. Once you do this, you can apply the mark ups and mark downs as you were doing.

OpenStudy (tkhunny):

"purchased snow shovels for $8 each. The original markup was 50% based on the selling price." Selling Price = $16.00 Markup = 16.00 - 8.00 = 8.00 Markup From Selling: 8.00/16.00 = 50% In December, the store had a sale and marked down the shovels by 20%. December Sale Price: 16.00 - 20%(16.00) = 16.00 - 3.20 = 12.80 By January 1, the sale was over and the shovels were marked up 15%. 12.80 + 15%(12.80) = 12.80 + 1.92 = 14.72 In March, the store held a clearance sale of all winter tools by having a final markdown of 60%. 14.72 - 60%(14.72) = 14.72 - 8.83 = 5.89 There are a couple of ambiguities in this path. Shown is one possible interpretation.